OpenStudy (anonymous):

Real Analysis : PLEASE please HELP : If g(x +y) = g(x) +g(y) for all x in R (real numbers). And g(x) > 0 for all x. and g(0) = 1 and g(x) >1 for some x in (0,a), a >0 then g is strictly increasing and continous on R.

OpenStudy (anonymous):

I meant g(x+y) = g(x)g(y)

OpenStudy (anonymous):

What is the question?

OpenStudy (kinggeorge):

Not sure if this helps, but the only function I know of that has the property g(x+y)=g(x)g(y) is exponentiation.

OpenStudy (kinggeorge):

Perhaps, since we know \(g(x+a)=g(x)\cdot g(a)\) if we let \(x=0\), and \(b\in (0, a)\) such that \(g(b)>1\) we have that \[g(0+b)=g(b)=g(0)\cdot g(b)=1\cdot g(b) > 1\]

OpenStudy (anonymous):

thanks so much I'm just thinking quick

OpenStudy (anonymous):

how do you show continuity? or that it's inreasing?

OpenStudy (anonymous):

increasing

OpenStudy (kinggeorge):

To show increasing, suppose \(g(b) \geq g(c)\) for some \(b<c\). Then, show that this implies a contradiction

OpenStudy (kinggeorge):

To show continuity, you need to show that for some \(\epsilon \in \mathbb{R}_{>0}\) there exists a \(\delta\in\mathbb{R}_{>0}\) such that \[|x-c|<\delta \implies |f(x)-f(c)|<\epsilon\]For all \(x\in\mathbb{R}\)

OpenStudy (kinggeorge):

And for all \(c\in\mathbb{R}\)

OpenStudy (anonymous):

do you know how to do the above fore this question in particular? I think i've shown that g must be increasing... I also know that if g is continous at x-0 then it is continous everywhere, i solved that

OpenStudy (anonymous):

for

OpenStudy (kinggeorge):

For it to be continuous at 0, use \(c=0\) in the above definition I gave you.

OpenStudy (kinggeorge):

So show that for every \(\epsilon\in\mathbb{R}_{>0}\) there exists a \(\delta\in\mathbb{R}_{>0}\) such that for all \(x\in\mathbb{R}\)\[|x|<\delta \implies |f(x)-1|<\epsilon\] I wrote my definition a little bit wonky up there.

OpenStudy (anonymous):

if g strictly increasing f(x)-f(c)=f(x-c) now (x-c)<delta f(x-c)<f(delta) ep=f(delta)??????????

OpenStudy (anonymous):

you guys are both saying very similar things to what i got in my solution, I'm just a bit woried about how to use the fact that g(x) >1 for all x in (0,a)

OpenStudy (anonymous):

and also the domains (-a,0) (0,a) for arbitrary a doesn't include the point 0